Suppose a hospital would like to estimate the proportion of patients who feel that physicians who care for them always communicated effectively when discussing their medical care. A pilot sample of 100 patients found that 70 reported that their physician communicated effectively. Determine the additional number of patients that need to be sampled to construct a 90​% confidence interval with a margin of error equal to 1​% to estimate this proportion.

Answer: The additional number of patients that need to be sampled is 5,582.

Above are the questions and answers. Please provide all Excel formulas (not calculations) that get to these answers.
Answer #1
here margin of error E = 0.01
for90% CI crtiical Z          = 1.645 # (from normsinv(0.95) function)
estimated prop.=p=70/100= 0.700
reqd. sample size n=p*(1-p)*(z/E)2=0.7*0.3*(1.645/0.01)^2 = 5683.00    

therefore additional number of patients that need to be sampled is =5683-100 =5583